Hi I need some help here please!!!!

Recall the earlier example of assessing the risk of loan defaults. Suppose

the bank’s top managers are divided on whether to adopt the scoring

system permanently. A number of top officers believe their intuitive

judgment about risks is superior to an “artificial” score. Accordingly, the

bank decides to test its judgment against the scoring system. The

managers will make their own designations of loans to the four

categories and see how well they can identify problem loans. Their track

record over the past year is shown in the table:

Category Performing Loan Defaulted Loan

zero risk 0.25 0.20

solid 0.30 0.25

uncertain 0.40 0.45

high risk 0.05 0.10

total 1 1

a. Predict the probability of default for each loan category. (Assume the overall default rate is 10 percent: Pr(default) .1.) b. How do these risk assessments, based on judgment and intuition, compare with the earlier predictions based on credit scores? Which seems to provide more valuable information? Explain.
Answer #1
a.
Category Performing Loan Defaulted Loan Probability of default(defaulted loan/Total loan) ie. Overall default rate 10%*individual probability
Zero Risk 0.25 0.2 0.2/(0.25+0.2)= 0.44 44.44% 4.44%
Solid 0.3 0.25 0.25/(0.3+0.25)= 0.45 45.45% 4.55%
Uncertain 0.4 0.45 0.45/(0.4+0.45)= 0.53 52.94% 5.29%
High risk 0.05 0.1 0.1/(0.05+0.1)= 0.67 66.67% 6.67%v